### v3

# Demographic-aware fine-grained classification of pediatric wrist fractures

Abstract
Wrist pathologies are frequently observed, particularly among children who constitute the majority of fracture cases. Computer vision presents a promising avenue, contingent upon the availability of extensive datasets, a notable challenge in medical imaging. Therefore, reliance solely on one modality, such as images, proves inadequate, especially in an era of diverse and plentiful data types. In this study, we employ a multifaceted approach to address the challenge of recognizing wrist pathologies using an extremely limited dataset. Initially, we approach the problem as a fine-grained recognition task. Secondly, we enhance network performance by fusing patient metadata with X-rays. Thirdly, we improve the performance further by utilizing weights trained on a separate fine-grained dataset. While metadata integration has been used in other medical domains, this is a novel application for wrist pathologies.
Index Terms—  Pediatric fracture classification, fracture detection, medical X-ray imaging, fine-grained classification, Patient metadata
1 Introduction
Hand injuries are among the most frequent presentations in emergency departments, with wrist fractures accounting for up to 75% of cases [1]. Given the wrist’s complex anatomy and critical role in hand function, accurate diagnosis is essential to prevent long-term impairment. Recent advances in computer vision have shown promise in detecting trauma-related abnormalities on X-rays; however, such methods typically require large, well-annotated datasets, which remain scarce in medical imaging. To overcome this limitation, multimodal learning approaches that combine patient metadata with imaging have been explored in domains such as dermatology and oncology, showing consistent improvements in diagnostic accuracy [2, 3, 4, 5, 6].
A consistent trend in these studies is the predictive value of demographic metadata, particularly age and sex. Thomas et al. [6] demonstrated accuracy gains by fusing age, sex, and anatomical site with dermoscopy images, while Li et al. [5] reported 89.1% accuracy on ISIC 2018 using PNASNet with metadata fusion. Pacheco et al. [7, 8] showed balanced accuracy improvements and introduced specialized fusion modules, and Ningrum et al. [2] achieved 92.3% melanoma classification by coupling ANN-based metadata encoders with CNNs. More recently, Cai et al. [4] applied ViTs with metadata encoders, reaching 93.8% accuracy. Collectively, these works establish metadata as a critical complement to imaging in medical AI.
Despite this progress, metadata integration has not been investigated for wrist pathology classification. This omission is critical because wrist anatomy and pathology presentation are strongly age- and sex-dependent. Carpal ossification occurs in well-defined stages, the distal radial and ulnar physes close at different ages for boys and girls [9, 10], and common pediatric fracture types (e.g., buckle, greenstick) differ from adult patterns [11, 12]. Without such demographic context, radiographic findings may be misinterpreted as abnormal when they are not, or vice versa.
To address this gap, we extend metadata integration to musculoskeletal imaging, framing wrist abnormality recognition as a fine-grained visual recognition (FGVR) task. Unlike prior multimodal efforts that primarily used conventional CNNs, we fuse demographic metadata with radiographs to capture subtle, age- and sex-specific morphological cues that image-only models often overlook. Building on our prior work image-only fine-grained recognition work [13, 14], we propose a joint framework for pediatric wrist pathology analysis and hypothesize that it will outperform image-only and CNN-based baselines.
Fig. 1: The architecture pipeline integrates visual, class, and meta tokens within a transformer framework.
2 Methods
2.1 Dataset Curation & Patient Metadata
We curated two datasets from GRAZPEDWRI-DX [15]. (1) A “limited” set of three wrist pathologies (fracture, bone anomaly, soft tissue), excluding the “metal” class, following our prior work [13, 14]. (2) A binary “Fracture vs. No Fracture” dataset using all images, with the non-fracture class augmented for balance. Final splits are shown in Table 1. The cohort included 6,091 patients (2,688 female, 3,402 male, ages 0.2–19 years, mean 10.9). Patient metadata is derived from the filename structure of each image. Fig. 2 shows distributions: fractures are more common in males, while anomalies concentrate in older groups.
Table 1: Instances in training, validation, and test sets.
Three Wrist Pathologies (Limited set) Fracture vs. No Fracture (Entire set)
Class Training Validation Test Class Training Validation Test
Boneanomaly 392 98 119 Fracture 11854 1357 338
Fracture 400 100 120 No Fracture 10164 1357 338
Soft-tissue 376 94 115
Total 1168 292 354 Total 22018 2714 676
Fig. 2: Bar and box plots for gender and age, respectively, for both curated datasets.
2.2 Architectural Details
This work utilizes a hybrid architecture that integrates both visual data and meta-information, drawing inspiration from the architecture called “MetaFormer” [16].
Visual Backbone. As depicted in Fig. 1, the architecture is organized into five stages. The first stage (S0) begins with a 3-layer convolutional stem that processes raw wrist images, introducing locality and inductive bias for edges and textures. Stages S1 and S2 consist of MBConv blocks enhanced with squeeze-and-excitation, which increase representational power by modeling inter-channel dependencies. These stages progressively downsample the feature maps while expanding channel dimensionality, ensuring the network can capture both low-level bone structures and mid-level soft tissue patterns. The final two stages, S3 and S4, transition from convolution to transformer-based blocks with relative positional encoding. This design allows the model to capture global dependencies, which are crucial for subtle pathological differences that span larger contexts (e.g., comparing radius vs. ulna contours or detecting diffuse soft tissue swelling).
Downsampling at the beginning of each stage reduces resolution while maintaining contextual continuity. In S3 and S4, overlapping patch embedding is employed via strided convolution with zero-padding [17], which provides partially redundant tokens that preserve boundary details, important when localizing hairline fractures or subtle cortical irregularities. Table 2 outlines the configuration details of the three variants of the architecture used in our experiments.
Table 2: Configuration of the architecture series. B represents the number of blocks, and H denotes the hidden dimensions for each stage.
Stage FG-0 FG-1 FG-2
S0 B=3 H=64 B=3 H=64 B=3 H=128
S1 B=2 H=96 B=2 H=96 B=2 H=128
S2 B=3 H=192 B=6 H=192 B=6 H=256
S3 B=5 H=384 B=14 H=384 B=14 H=512
S4 B=2 H=768 B=3 H=768 B=3 H=1024
Transformer Fusion. Standard self-attention lacks spatial awareness; therefore, we incorporate a relative position bias. Specifically, a bias matrix $B\in\mathbb{R}^{((M^{2}+N)\times(M^{2}+N))}$ is added during attention computation:
$\textit{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d}}+B\right)V$ (1)
Here, $Q$, $K$, and $V$ are query, key, and value matrices of size $\mathbb{R}^{((M^{2}+N)\times d)}$. $M^{2}$ corresponds to the number of image patches and $N$ is the count of additional tokens (class and meta). This positional encoding ensures that the model attends appropriately to the relative arrangement of wrist bones, where spatial context is essential for distinguishing true pathology from normal variants.
The Relative Transformer Block is composed of multi-head self-attention (MSA), layer normalization (LN), and a multi-layer perceptron (MLP). It operates on a token sequence $z_{0}$ that combines visual, meta, and class tokens:
$z_{0}$ $=[x_{\text{class}};x^{1}_{\text{meta}},\ldots,x^{n-1}_{\text{meta}};x^{1}_{\text{vision}},\ldots,x^{m}_{\text{vision}}]$ (2)
$z^{\prime}_{i}$ $=\text{MSA}(\text{LN}(z_{i-1}))+z_{i-1}$
$z_{i}$ $=\text{MLP}(\text{LN}(z^{\prime}_{i}))+z^{\prime}_{i}$
This formulation enables the network to learn correlations not only within visual patches but also across modalities. For example, the model can attend jointly to a distal radius patch and the token representing a 70-year-old female, thereby weighting its decision toward fracture likelihood based on both visual evidence and prior epidemiology.
Multi-Scale Aggregation. Stages S3 and S4 output separate class tokens, $z^{1}_{\text{class}}$ and $z^{2}_{\text{class}}$, which encode fused visual and meta representations at different scales. Since dimensions differ, $z^{1}_{\text{class}}$ is projected using an MLP to align with $z^{2}_{\text{class}}$. These are then concatenated and passed through a 1D convolution followed by normalization:
$\hat{z}^{\text{1}}_{\text{class}}$ $=\text{MLP}(\text{LN}(z^{\text{1}}_{\text{class}}))$ (3)
$z_{\text{class}}$ $=\text{Conv1d}(\text{Concat}(\hat{z}^{\text{1}}_{\text{class}},z^{\text{2}}_{\text{class}}))$
$y$ $=\text{LN}(z_{\text{class}})$
The final output $y$ aggregates both global image evidence and auxiliary patient priors, which is especially important in fine-grained medical classification.
Meta-information Integration. When external attributes are available, they are embedded into the same feature space as vision tokens. In our case, age and gender are encoded as low-dimensional vectors and passed through a non-linear fully connected embedding function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$. This ensures that demographic priors, e.g., higher fracture prevalence in elderly females, are directly fused with image tokens. Importantly, meta-information is gradually incorporated during training to prevent early over-reliance on attributes, allowing the visual encoder to remain robust.
To maintain the model’s ability to learn important visual features, especially when metadata is less informative or potentially overfitting, a regularization strategy is applied during training that progressively masks parts of the meta-information. This balances the reliance on both vision and auxiliary data.
2.3 Experimental Settings
A range of well-established models has been incorporated in our study alongside our fine-grained metadata-aware approach for comparison. All these baseline models were pre-trained on ImageNet [18]. These include ViT [19], VGG16 [20], GoogleNet [21], EfficientNetV2 [22], AlexNet [23], DenseNet201 [24], InceptionV3 [25], and ResNet50 [26]. All models underwent training for 100 epochs using an image resolution of 224x224 and a batch size of 32. Adam optimizer was utilized with a learning rate $\eta=5\times 10^{-3}$. For MetaFormer, however, the base learning rate was $\eta=1\times 10^{-4}$, warmup learning rate $\alpha=5\times 10^{-7}$, minimum learning rate was $\eta_{m}=1\times 10^{-5}$, and weight decay was $\wp=5\times 10^{-2}$.
3 Results & Discussion
Table 3: Performance evaluation of different baseline neural networks and the MetaFormer on only image data.
Model Test Accuracy (Vision Only)
ViT 62.71%
VGG16 63.28%
GoogleNet 70.34%
EfficientNetV2 72.32%
AlexNet 74.29%
DenseNet201 76.84%
InceptionV3 77.44%
ResNet50 77.68%
MetaFormer FGVR (Base) 79.40%
Table 4: Accuracy of different MetaFormer configurations on only image data, integrating metadata through early fusion, and integrating metadata through late fusion.
Model Configuration Vision Only Early Fusion Late Fusion
FG-0 79.4% 79.1% 80.5%
FG-1 78.5% 80.5% 79.1%
FG-2 77.4% 78.5% 79.4%
FG-2-inat (Ours) 79.9% 81.1% 81.4%
Table 5: Evaluation of our configuration of MetaFormer (FG-2-inat) using sensitivity, specificity, and precision.
Configuration Cls Sensitivity Specificity Precision
Vision Only 0 62.2% 98.7% 96.1%
1 100% 82.9% 75%
2 77.4% 88.3% 76.1%
Vision+Meta 0 66.4% 98.7% 96.3%
1 100% 82.9% 75%
2 77.4% 90.4% 79.5%
3.1 Baseline Comparison
Table 3 shows that MetaFormer (79.4%) outperforms widely used CNNs such as ResNet50 (77.7%) and DenseNet201 (76.8%). This suggests that transformer-based architectures with fine-grained design are better suited to capture subtle morphological cues in pediatric wrist radiographs, where differences between normal anatomy and pathology can be minimal, giving credence to our previous works [13, 14].
3.2 Metadata Integration & Pretraining
Integrating demographic metadata (age, sex) consistently improved performance (Table 4). Both early and late fusion surpassed vision-only models, with late fusion achieving the highest gains in two out of three configurations. This finding aligns with the intuition that age and sex provide essential developmental context, particularly relevant in pediatrics, where normal ossification patterns vary across individuals. Attribute-level analysis (Table 6) further indicates that both age and sex independently contribute to performance, but their combination yields the largest benefit, supporting the clinical relevance of demographic context.
The best configuration “FG-2-inat” with fine-grained pretraining reached 81.4%, a 1.5% improvement over image-only training. Fine-grained pretraining on iNaturalist [27] improved generalization beyond ImageNet pretraining, reflecting the value of transferring from domains where subtle inter-class variations dominate.
On the larger “Fracture vs. No Fracture” dataset (Table 7), metadata integration produced a 10% absolute gain, underscoring the robustness of multimodal learning even when the task is coarser. These results suggest that metadata not only aids fine-grained wrist pathology recognition but also enhances scalability to larger clinical datasets.
Fig. 3: Heatmaps generated from the final convolutional layer in stage 2 (S2) of our configuration fine-grained-FG-2 on a limited dataset.
3.3 Interpretability
GradCAM heatmaps (Fig. 3) demonstrate that MetaFormer attends to clinically relevant wrist regions, particularly around fracture lines and anomalous bone structures. While most heatmaps align well with pathology, some exhibit diffuse attention, suggesting occasional reliance on spurious features. This indicates that while metadata integration improves classification, further refinement is needed for precise localization. Nevertheless, the alignment of salient regions with clinical expectations strengthens confidence in the model’s decision process.
Table 6: Table showing accuracies obtained by our configuration of MetaFormer on individual metadata attributes and when both attributes are considered.
Attribute Early Fusion Late Fusion
Vision + Age 80.5% 79.4%
Vision + Gender 79.9% 80.8%
Vision + Age + Gender 81.1% 81.4%
Table 7: Performance evaluation on significantly larger (fracture vs. no fracture) dataset.
FG-2-inat (Ours) Test Accuracy
Vision Only 50.1%
Vision + Meta 60.4%
4 Conclusion
In this study, we demonstrated that integrating patient demographic metadata with image data enhances fine-grained recognition of pediatric wrist pathologies. Overall, our experiments demonstrate three key findings: (1) Fine-grained transformer architectures outperform conventional CNNs for pediatric wrist pathology. (2) Demographic metadata (age, sex) significantly improves recognition, with late fusion being most effective. (3) Fine-grained pretraining enhances generalization, and metadata boosts performance across both limited and large-scale datasets. These results emphasize that integrating developmental context into vision models is not only beneficial but essential for robust pediatric musculoskeletal diagnosis.
5 Acknowledgement
This work was supported by the Curricula Development and Capacity Building in Applied Computer Science for Pakistani Higher Education Institutions (CONNECT) Project NORPART-2021/10502, funded by the Norwegian Directorate for Higher Education and Skills (DIKU).